ONE HUNDRED VOLUMES OF PUBLICATIONS DE L’INSTITUT MATHÉMATIQUE

Publications de l’Institut Mathématique is being published continuously since 1947, and this issue is the 100th volume in the series. Here we celebrate long life and successful tradition of the journal. The history of Publications is rich, interesting and it is tightly connected with the history of mathematics in Serbia and the history and development of the Mathematical Institute of the Serbian Academy of Sciences and Arts.

In the mid-twenties of the XX century, a strong group of mathematicians was working at the Belgrade University. The older generation was represented by Bogdan Gavrilović, Mihailo Petrović and Milutin Milanković, while the younger mathematicians were Tadija Pejović, Jovan Karamata, Radivoj Kašanin, Miloš Radojčić and a few others. Several mathematicians from Russia who had to flee their country because of the October revolution were also present at the time. Anton Bilimović came from Imperator Novorossian University in Odessa, while Nikola Saltikov was from Kharkov State University. They were distinguished mathematicians with a very good mathematical background. For example, the world famous mathematicians Vladimir Andreevich Steklov and Aleksandr Mikhailovich Lyapunov were the advisors for the doctoral dissertation of Nikola Saltikov.

Following the initiative of Anton Bilimović, the Club of Mathematicians of the Belgrade University was soon founded. The Club did not have any specific rules, except that it gathered once a month in order for the members to present their current work and papers. At that time there was no purely mathematical journals published in Serbia, so Serbian mathematicians published their papers in foreign journals, as well as in the general journal Voice of the Serbian Royal Academy and in “Rad” (Work, in English) of the Yugoslav Academy of Sciences and Arts (now the Croatian Academy of Sciences). However, the rule of the Serbian Academy was that the papers need to be written in Serbian, hence the papers printed in the Voice in most cases remained unknown to the international mathematical community. The actual need for a Belgrade-based mathematics journal was present. Once again thanks to the initiative of Anton Bilimović, the first Serbian mathematics journal was established in 1932, under the name Publications Mathématiques de l’Université de Belgrade. Two Belgrade University trusts “Luka Ćelović-Trebinjac” and “Pavle and Katarina Kurtović” helped its foundation. Papers were written in French and German. Not only that the leading Serbian mathematicians published in Publications at the time, but the list of authors included other Yugoslav and many recognized world mathematicians, most prominent of them being Wilhelm Blaschke, Ellie Cartan, Henri Lebesgue, Paul Montel, Josip Plemelj, Waclaw Sierpinski, etc. Many fine papers appeared there, some of them of capital importance and significantly influencing several areas of mathematics. For example, Ðuro Kurepa published in extenso his doctoral dissertation Ensembles ordonnées et ramifiés (vol. IV, 1935), where the set-theoretical notion of the tree, as well as many other related notions, was introduced. Jovan Karamata, in his paper Sur une inégalité relative aux fonctions convexes (vol. I, 1932) proved his remarkable inequality, which is a generalization of the Jensen’s inequality. Seven volumes were published until World War II, while the eighth volume was lost in the printing shop during the German bombing of Belgrade, that occurred on April 6, 1941. The journal was not published during the period 1941–1947.

Immediately after the World War II, the Academy announced the establishment of the Mathematical Institute on its IX gathering held on October 18, 1945. The final decision was made on the first assembly of the Academy on April 25, 1946, and this date is taken as the day of the foundation of the Mathematical Institute. Soon after that, in 1947, the publication of the journal restarted under the new name Publications de l’Institut Mathématique de l’ Académie Serbe des Sciences et des Arts. The journal started with the new numeration, beginning from No 1. Therefore, this issue of the Publications is in fact VII+100 (107th) volume, if we include the volumes published before World War II. In 1961, the Mathematical Institute became independent from the Academy and then the journal changed its name to Publications de l’Institut Mathématique and started the new numeration (nouvelle série), but the old one was kept as well. The Institute returned later (in 1995) the previous name, but the name of the journal was not changed.

Up until now, authors from more than 30 countries published in total 2112 articles in Publications. This makes the journal truly international. In the beginning, the scope of the journal was broader, including not only mathematics, but also mechanics and astronomy. Most prominent Serbian and Yugoslav scientists in these fields, including Slobodan Aljančić, Jovan Karamata, Ðuro Kurepa, Svetozar Kurepa, Sibe Mardešić, Milutin Milanković, Miodrag Tomić, Ivan Vidav, and others, continued to publish their papers in this journal. Some of the wellknown mathematicians published in Publications as well: János Aczél, Nicholas H. Bingham, Paul Erdős, Laurens De Haan, Marek Kuczma, Saharon Shelah, Eugene Seneta and others. Currently, most papers published in the journal are in English, but papers written in Russian, French, and German are accepted as well. The Library of the Mathematical Institute receives the copies of more than 200 journals as a result of extensive exchange of scientific publications with different institutions from all over the world.

The policy of the Mathematical Institute and the Editorial Board of Publications is to provide a journal of high standards. Publications invites submission of original scholarly work in mathematics and its applications. The journal features selected high quality articles that represent the broad spectrum of research in contemporary mathematics. It is particularly intended to promote those areas of mathematics which reflect the research interest of Serbian mathematicians. Occasionally, Publications publishes volumes dedicated to selected topics of wider interest. Besides research papers, articles can be purely expository or historical reviews.

Two volumes of the journal are published per year, in June and November. The complete retrodigitization of the journal was finished in 2006, and this electronic version is now available on the Web page of the Mathematical Institute. The articles published after 1980 are prepared in TEX, and they are ready for indexing by Internet search engines. The Mathematical institute stands by the idea that the mathematical results should be freely shared within the scientific and mathematical community, and as a result downloading articles from Publications is free of charge.

Some volumes of the journal were dedicated to distinguished Serbian and Yugoslav mathematicians: Mihailo Petrović (vol. VI–VII, 1938), Josip Plemelj (vol. 4(18), 1964), Slobodan Aljančić (vol. 58(72), 1995), Ðuro Kurepa (vol. 57(71), 1995), Jovan Karamata (vol. 48(62), 1990 and vol. 71(85), 2002) and Aleksandar Kron (vol. 82(96), 2007).

In the sequel, we list all members of the Editorial Board of the journal. Written records do not exist for the period 1932–1941. However, we can indirectly conclude that Anton Bilimović was the main organizer, and that the other members of the Belgrade Club of Mathematicians were involved in the preparation of the journal, in particular Mihailo Petrović, Milutin Milanković, Jovan Karamata and Bogdan Gavrilović. Until 1984 the Editor-in-Chief was not explicitly named, but members elected him at the first meeting of the new Board. From 1984 on, the Scientific Council of the Mathematical Institute appoints the Editor-in-Chief. In 1994 the Editorial Board was reorganized and since then it consists from editors and the Editorial Board. While the editors are in charge for handling papers in particular areas of mathematics, the Editorial Board has mainly the advising role and it consists of the representatives of the Serbian Academy of Sciences and Arts, foreign members and Serbian mathematicians in Diaspora.

Members of the Editorial Board of Publications

1947–1952: A. Bilimović, V. Mišković, T. Anđelić, later R. Kašanin, J. Karamata, and V. Avakumović joined the Board. 1953: R. Kašanin (Editor-in-Chief), A. Bilimović, V. Mišković, J. Hlitčiev, T. Anđelić. 1953–1955: R. Kašanin (Editor-in-Chief), A. Bilimović, V. Avakumović. 1956: R. Kašanin (Editor-in-Chief), A. Bilimović, T. Anđelić. 1957–1958: R. Kašanin (Editor-in-Chief), V. Avakumović, M. Tomić. 1959–1960: R. Kašanin (Editor-in-Chief), M. Tomić, S. Aljančić. 1961–1964: T. Pejović (Editor-in-Chief), D. Mitrinović, D. Marković, T. Anđelić, B. Rašajski. 1965–1966: T. Pejović (Editor-in-Chief), D. Mitrinović, T. Anđelić, B. Rašajski, V. Brčić.

1967–1969: T. Pejović (Editor-in-Chief), D. Mitrinović, T. Anđelić, B. Rašajski, S. Pivko. 1970–1971: T. Anđelić (Editor-in-Chief), M. Barjaktarević, S. Kurepa, D. Mitrinović, T. Pejović, S. Pivko, B. Popov, B. Rašajski, I. Vidav. 1972–1983: S. Aljančić (Editor-in-Chief), M. Barjaktarević, B. Popov, S. Prešić, M. Prvanović, B. Rašajsky, I. Vidav, V. Vujičić. D. Cvetković, G. Čupona, A. Ivić, Ð. Kurepa, S. Kurepa, V. Marić, M. Marjanović, Ž. Mijajlović, S. Prešić, M. Prvanović, I. Vidav. 1994–2005: Editors: M. Prvanović (Editor-in-Chief), A. Ivić, M. Jevtić, Ž. Mijajlović, G. Milovanović, S. Pilipović, S. K. Simić, R. Živaljević. Editorial Board: P. Ahern (USA) (only in 1994), D. Aranđelović (Serbia), A. Björner (Sweden), M. Borowiecki (Poland), D. Cvetković (Serbia), A. Ivić (Serbia), S. Janković (Serbia, since 2003), Z. Ivković (Serbia), V. Kudryavcev (Russia), V. Marić (Serbia), M. Oberguggenberger (Austria), S. Prešić (Serbia), B. Rosenfeld (USA), Yu. Solovyev (Russia, since 1995), G. Tenenbaum (France), S. Todorčević (Serbia) 2006– : Editors: Ž. Mijajlović (Editor-in-Chief), V. Dragović, A. Ivić, S. Janković, M. Jevtić, M. Kurilić, D. Milinković, G. Milovanović, S. Pilipović, S. K. Simić, R. Živaljević. Editorial Board: A. Björner (Sweden), M. Borowiecki (Poland), D. Cvetković (Serbia), V. Kudryavcev (Russia), E. Omey (Belgium), M. Oberguggenberger (Austria), B. Rosenfeld (USA), G. Tenenbaum (France), S. Todorčević (Serbia)

Secretaries: V. Popović till 1980, M. Čavčić (1980–1982), T. Ostrogorski (1982– 2005), B. Gajić since 2006. Technical editors: M. Čavčić till 1982, since then D. Blagojević.

The following sources were used in preparation of this text: 1. M. Čavčić, Review of publications of Mathematical Institute 1946–1986 (in Serbian), Math. Inst. Belgrade, 1986. 2. Life and Works of Serbian Scientists, vol. 1–10, ed. M. Sarić, V. Ðorđević, Serb. Acad. Sci. Arts, published 1996–2009. 3. Virtual Library of the Faculty of Mathematics University of Belgrade (http://elib.matf.bg.ac.rs: 8080/virlib). 4. Publications de l’Institut Mathématique home page (http:// publications.mi.sanu.ac.rs).

Žarko Mijajlović